OpenStudy (anonymous):

If Gunnar gives me a cookie, I will have twice as many as he. If I give Gunnar a cookie, we will have the same number of cookies. How many cookies do I have? How many cookies does Gunnar have? Explain in detail how you found your answer using words, number, and/or pictures.

OpenStudy (amistre64):

j+1 = 2g j-1 = g j+1 = 2(j-1) j+1 = 2j -2 3 = j 2 = g